# Distinctive Feature Codec: Adaptive Segmentation for Efficient Speech Representation

### v1

@@ Abstract
The tokenization of speech with neural speech codec models is a crucial aspect of AI systems designed for speech understanding and generation. While text-based systems naturally benefit from token boundaries between discrete symbols, tokenizing continuous speech signals is more complex due to the unpredictable timing of important acoustic variations. Most current neural speech codecs typically address this by using uniform processing at fixed time intervals, which overlooks the varying information density inherent in speech. In this paper, we introduce a distinctive feature-based approach that dynamically allocates tokens based on the perceptual significance of speech content. By learning to identify and prioritize distinctive regions in speech signals, our approach achieves a significantly more efficient speech representation compared with conventional frame-based methods. This work marks the first successful extension of traditional signal processing-based distinctive features into deep learning frameworks. Through rigorous experimentation, we demonstrate the effectiveness of our approach and provide theoretical insights into how aligning segment boundaries with natural acoustic transitions improves codebook utilization. Additionally, we enhance tokenization stability by developing a Group-wise Scalar Quantization approach for variable-length segments. Our distinctive feature-based approach offers a promising alternative to conventional frame-based processing and advances interpretable representation learning in the modern deep learning speech processing framework.
1 Introduction
Figure 1: Frame processing vs. Distinctive processing. Distinctive processing identifies natural acoustic boundaries for variable-length segments.
The remarkable success of large language models (LLMs) in understanding and generating text [1, 2, 3] has inspired researchers to develop similar architectures for speech processing. This expansion towards speech is motivated by the rich information encoded in speech signals beyond mere linguistic content - including speaker identity, emotion, prosody, and acoustic environments [4, 5, 6, 7] - which could enable AI systems to better understand and interact with humans [8, 9]. Indeed, recent speech language models have demonstrated impressive performance across various tasks, from speech recognition to speech generation [10, 11, 12]. A fundamental component in building these speech LLMs is the tokenization of speech signals - the process of converting continuous audio into discrete tokens that can be effectively processed by neural networks. However, unlike text where natural boundaries exist between words or subword units [13, 14, 15], speech signals present unique challenges for tokenization due to their continuous nature and the complex interplay of acoustic features. Most current approaches primarily rely on frame-based processing with fixed time intervals [16, 17, 18], leading to significant redundancy in the resulting representations. This redundancy manifests as longer sequences of speech tokens compared with their text counterparts, making speech LLMs substantially more challenging to train and deploy. Moreover, frame-based processing often fails to capture the inherent structure of speech, where important acoustic variations can occur at multiple time scales [19] and linguistic boundaries may not align with fixed frame intervals [20]. These limitations have motivated researchers to seek alternatives to frame-based processing for more efficient speech representation  [^1] Audio Sample site:https://sites.google.com/view/codec-audio-sample/home
Code:https://github.com/Tonyyouyou/Distinctive-Codec
The key to efficient speech representation lies in identifying meaningful units within continuous signals. While text naturally segments into distinct units, speech requires a principled approach to find its fundamental elements. The linguistic theory of distinctive features provides an insightful perspective: speech can be naturally segmented at points where acoustic characteristics are most differentiated [21, 22, 19, 23]. These theories suggest that not all parts of speech carry equal information - rather, the most informative regions occur where acoustic properties undergo significant changes, marking natural boundaries between perceptually distinct segments [24, 25]. This perspective offers a compelling alternative to frame-based processing: instead of arbitrarily dividing speech into fixed-length segments, we can identify boundaries where speech characteristics are most differentiated. However, while frame-based processing has seamlessly integrated with deep learning through operations like convolution, which naturally processes fixed-interval frames [26, 27], implementing distinctive feature analysis in neural networks has remained challenging. The irregular nature of distinctive features - their varying lengths and dynamic characteristics - conflicts with the regular, grid-like computations that have dominated modern deep learning architectures [28, 29]. This technical mismatch, rather than any fundamental limitation of the theory itself, has largely confined distinctive feature analysis to traditional signal processing approaches [20, 30, 23].
In this paper, we address the challenge of efficiently tokenizing continuous speech signals by drawing inspiration from the linguistic concept of distinctive features. Using the encoder-decoder based codec framework [31, 32] with SEANet [33] which has been used in many speech codec models [16, 17, 18] as our experimental platform, we demonstrate that identifying and emphasizing perceptually distinct regions leads to more efficient speech tokenization compared with fixed-interval processing. Our systematic evaluation in the codec setting shows that this approach not only achieves higher compression efficiency but also better preserves critical acoustic properties across various speech characteristics. Through theoretical analysis, we provide insights into why distinctive feature-based processing succeeds in speech codec models, establishing a foundation for future codec development. Additionally, our investigation reveals and addresses a fundamental limitation in Finite-State Quantization (FSQ) [34] when operating at high downsampling rates. By analyzing the underlying causes of instability, we developed an improved quantization scheme that maintains robust performance at high downsampling rates. Our work validates distinctive features as a promising direction for codec design, offering new perspectives on efficient speech representation.
2 Preliminary
Distinctive Features
@@ 2 Preliminary
first proposed in linguistic theory [21, 22, 19, 23], characterize speech by identifying regions with acoustically distinctive properties that help differentiate speech segments from one another. As illustrated in Fig.1, this approach fundamentally differs from conventional frame-based processing: while frame-based methods uniformly segment speech signals into fixed-length overlapping windows, distinctive feature analysis identifies boundaries where acoustic characteristics undergo significant changes. This approach has proven valuable in early automatic speech recognition systems [22, 19, 35] and has been successfully applied to various healthcare applications [36, 37, 38, 39, 40]. However, despite its theoretical advantages, the development of distinctive feature-based methods has faced significant limitations in the deep learning era. Traditional implementations of distinctive features heavily rely on linguistic expertise and hand-crafted rules [22, 23], leading to limited training data and difficulties in scaling across different acoustic conditions and languages. Additionally, while frame-level processing naturally aligns with convolutional neural networks and enables efficient batch processing, the variable-length nature of distinctive features poses challenges for modern deep learning architectures. The success of frame-level processing in various deep learning systems [26, 27] has led to a rich ecosystem of pre-trained models and established practices, making it the predominant choice for modern speech processing systems despite its inherent inefficiencies. This technical mismatch, combined with the limited availability of labeled data for distinctive feature analysis, has constrained its adoption in contemporary deep learning approaches.
Speech Codecs
aim to compress speech signals into discrete tokens while preserving essential acoustic and linguistic information [41, 42]. These discrete representations can be used for efficient storage, transmission, and as input for downstream tasks such as large language models [11, 12, 10]. Current speech codec approaches can be broadly categorized into two paradigms: end-to-end trained models like SpeechTokenizer that directly learn the discrete representations [17, 16, 18], and two-stage approaches that first extract speech discrete features [11, 12] using self-supervised models [27, 26] followed by generative modeling techniques such as flow matching [43] or diffusion models [44]. The latter category, particularly approaches utilizing flow matching-based models, has demonstrated impressive capabilities in achieving high compression rates, reducing the number of tokens per second to remarkably low levels [45, 46, 47]. However, multiple studies have shown that these highly compressed representations often underperform in downstream LLM tasks when operating at low frame rates [48, 49]. We hypothesize that this performance degradation occurs because diffusion-based models’ strong generative capabilities are actually ’guessing’ missing speech information, resulting in representations that may not contain sufficient original speech information. We provide some exploratory experiments supporting this hypothesis in Appendix A. Given these considerations, our experimental investigation primarily focuses on end-to-end trained codecs, which we believe provide a more direct and transparent approach to learning speech representations.
3 Distinctive Codec Framework
@@
Our approach consists of two key stages: First, a lightweight boundary detector is trained to identify perceptually significant transitions in speech signals. These detected boundaries then guide the codec model to adaptively merge or separate speech segments, leading to more efficient tokenization that aligns with the natural structure of speech.
Distinctive Features Detector
@@
The core idea of distinctive features lies in identifying regions where speech segments exhibit maximal acoustic contrast with their neighbors. This naturally aligns with the objective of contrastive learning, which aims to learn representations by maximizing the differences between distinct samples while minimizing differences between similar ones [50, 51, 52, 53]. We leverage this connection to design a self-supervised boundary detector that learns to identify distinctive features without requiring phoneme-level annotations.
@@
Figure 2: The Overview Framework of Distinctive Codec, DF represents Distinctive Features
Distinctive Codec
To evaluate our distinctive feature-based approach, we build our codec model on top of the SpeechTokenizer framework [18], which leverages the SEANet architecture [33] for encoder-decoder operations. The key innovation of our approach lies in how we process the encoded representations based on the distinctive features detected by our boundary detector. Given input speech $\mathbf{x}\in\mathbb{R}^{1\times L}$, the encoder $E_{\theta}$ first maps it to a latent representation:
@@
where $D$ is the feature dimension and $T=L/r$ represents the temporal dimension after downsampling with ratio $r$. Traditional frame-based codecs would typically process this representation uniformly across time. In contrast, our distinctive codec first identifies segment boundaries $\mathcal{B}={b_{1},b_{2},\ldots,b_{M}}$ using the boundary detector described in Section 3. With these boundaries, we partition the feature sequence into variable-length segments:
@@
where $\mathbf{s}i=\mathbf{e}[:,b_{i-1}:b_{i}]$ represents the $i$-th segment ($b_{0}=0$ and $b_{M+1}=T$ for notational convenience). For each segment, we apply a Distinctive Feature encoder (DFE) to compress the variable-length representation into a fixed-length embedding:
@@
$\hat{\mathbf{x}}=D_{\phi}(\hat{\mathbf{e}})$ (8)
Group-wise Scalar Quantization
Finite Scalar Quantization (FSQ) [34] has emerged as an effective approach for discrete representation learning due to its computational efficiency and strong performance across various tasks [12, 54, 55]. Unlike vector quantizers that require nearest neighbor search in high-dimensional spaces, FSQ directly quantizes each dimension of the latent representation independently, significantly reducing computational complexity.
@@
However, during our experiments with distinctive feature-based tokenization, we discovered that standard FSQ becomes unstable when operating at high downsampling rates. This instability manifests as training divergence and poor reconstruction quality, particularly when compressing longer segments into single tokens. We hypothesize that this issue stems from the increased difficulty of directly quantizing high-dimensional features with varying temporal characteristics.
To address this challenge, we propose a Group-wise Scalar Quantization (GSQ) approach that enhances stability while maintaining the computational benefits of FSQ. The core idea is to decompose the high-dimensional quantization problem into multiple lower-dimensional sub-problems that can be solved more effectively.
Given a compressed segment representation $\mathbf{z}_{i}\in\mathbb{R}^{H\times 1}$, we first divide it into $G$ groups, each with dimension $H_{g}=H/G$:
$\mathbf{z}_{i}=[\mathbf{z}_{i}^{1},\mathbf{z}_{i}^{2},...,\mathbf{z}_{i}^{G}]$ (9)
Each group is then processed independently through its own specialized quantization path. This group-wise processing can be formalized as:
$\hat{\mathbf{z}}_{i}^{g}=Q_{g}(\mathbf{z}_{i}^{g})$ (10)
where $Q_{g}$ represents the group-specific quantization function that may include dimensionality reduction, scalar quantization, and dimensionality restoration. Each group’s quantizer preserves the most relevant information for that specific feature subspace, allowing for more nuanced representation of complex acoustic patterns.
@@
$\hat{\mathbf{z}}_{i}=[\hat{\mathbf{z}}_{i}^{1},\hat{\mathbf{z}}_{i}^{2},...,\hat{\mathbf{z}}_{i}^{G}]$ (11)
This group-wise approach is particularly beneficial when quantizing variable-length segments in our distinctive feature-based codec. The diverse acoustic characteristics present in different segments require adaptable quantization strategies. By processing features in smaller groups with dedicated quantizers, we reduce the complexity of the quantization problem through lower-dimensional operations. A notable implementation in our codec employs a many-to-one mapping architecture where high-dimensional feature groups are first projected to single scalar values through learned compression matrices. These scalar values undergo quantization via FSQ, and are subsequently projected back to their original dimensionality through expansion matrices:
$p_{i}^{g}=\mathbf{W}_{g}\mathbf{z}_{i}^{g},\quad\hat{p}_{i}^{g}=\text{FSQ}_{g}(p_{i}^{g}),\quad\hat{\mathbf{z}}_{i}^{g}=\mathbf{V}_{g}\hat{p}_{i}^{g}$ (12)
4 Experimental Results
Experimental Setup
We implemented our Distinctive Codec strictly(matched all codec settings, such as the model configuration, downsampling ratio) following the SpeechTokenizer framework [18] as our foundation because it effectively balances acoustic fidelity with semantic information while using an architecture widely adopted across modern speech codec research. This framework employs the SEANet-based encoder-decoder structure that has become a standard approach in numerous high-performance speech codecs, including SoundStream [17], EnCodec [56], and DAC [56]. Our model architecture employed the same SEANet-based structure as SpeechTokenizer, maintaining consistent hyper-parameters and loss functions to ensure a fair comparison that isolated the impact of our distinctive feature-based approach. Importantly, our main experiments did not include the HuBERT distillation component used in the original SpeechTokenizer, as we aimed to focus solely on the architectural comparison of frame-based versus distinctive feature-based processing. We did, however, explore the effect of distillation in Appendix D.1 to provide a complete analysis. For training, we used the LibriSpeech dataset [57], specifically its train set, because it represents a widely-used benchmark in speech processing research with sufficient diversity in speakers and content. We evaluated on 500 randomly selected samples from the test set, following the evaluation protocol established in SpeechTokenizer to maintain consistency with prior research and enable direct comparison of results. The Distinctive Features Detector uses a CNN architecture with 5 convolutional layers. For detailed model configurations, hyperparameters, and training procedures, please refer to Appendix B.
Evaluation Metrics
We evaluate our Distinctive Codec using metrics for reconstruction quality, intelligibility, and encoding efficiency following previous works [17, 16]. For reconstruction, we measure mel-spectral error, STFT distance, and PESQ (Perceptual Evaluation of Speech Quality) scores [58]. For intelligibility, we measure Word Error Rate (WER) using the Whisper en-medium model [59], following SpeechTokenizer [18], and STOI (Short-Time Objective Intelligibility) [60] to quantify how accurately the speech content is preserved. We also track encoding efficiency through Token Ratio (TKR) [61], representing tokens per second of 16 kHz audio, and Bits Per Second (BPS). The BPS calculation follows the approach in [55] which will consider the vocabulary size of quantization method. Detailed descriptions of these metrics are provided in Appendix C.
Main Result
Table 1: Performance comparison of different speech tokenization methods. Frame: frame number in one second; TKR: tokens per second [61]; BPS: bits per second [55]. RVQ refers to Residual Vector Quantizer [62]. FSQ refers to Finite Scalar Quantization [34]. Lower values are better for MEL Error, STFT, and WER; higher values are better for PESQ [58] and STOI [60]. All models follow identical training and architectural settings derived from SpeechTokenizer [18], ensuring fair comparison.
Model Quantize Method Frame TKR BPS Metrics

### v2

# Distinctive Feature Codec: An Adaptive Efficient Speech Representation for Depression Detection

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable success across diverse fields, establishing a powerful paradigm for complex information processing. This has inspired the integration of speech into LLM frameworks, often by tokenizing continuous audio via neural speech codecs, enabling powerful speech language models. However, this dominant tokenization strategy relies on uniform frame-based processing at fixed time intervals. This fixed-rate approach, while effective for linguistic content, destroys the temporal dynamics. These dynamics are not noise but are established as primary biomarkers in clinical applications such as depression detection. To address this gap, we introduce the Distinctive Feature Codec (DFC), an adaptive framework engineered to preserve this vital timing information. Drawing from linguistic theory, DFC abandons fixed-interval processing and instead learns to dynamically segment the signal at perceptually significant acoustic transitions. This generates variable-length tokens that efficiently encode the temporal structure. As a key contribution, this work is the first to integrate traditional distinctive features into a modern deep learning codec for a temporally sensitive task such as depression detection. We also introduce the Group-wise Scalar Quantization (GSQ) approach to stably quantize these variable-length segments. Our distinctive feature-based approach offers a promising alternative to conventional frame-based processing and advances interpretable representation learning in the modern deep learning speech depression detection framework.
I Introduction
The remarkable success of large language models (LLMs) in understanding and generating text [bai2023qwen, dubey2024llama, touvron2023llama] has inspired researchers to develop similar architectures for speech processing [defossez2024moshi, du2024cosyvoice, du2024cosyvoice2]. This expansion is motivated by the rich information encoded in speech signals beyond mere linguistic content, including speaker identity, emotion, and prosody [borsos2023audiolm, nguyen2023generative, lakhotia2021generative, zhang2023speechgpt]. A fundamental challenge in building these speech-aware models is the tokenization of continuous audio into representations that neural networks can process. Unlike text, which has natural boundaries [yu2023megabyte, the2024large, pagnoni2024byte], speech is continuous and complex. Consequently, most current approaches—whether for discrete tokenization [defossezhigh, zeghidour2021soundstream, zhang2024speechtokenizer] or self-supervised feature extraction [baevski2020wav2vec, hsu2021hubert]—primarily rely on frame-based processing with fixed time intervals.
While this fixed-rate processing is effective for tasks focused on linguistic content, such as speech recognition, it presents a fundamental limitation: it overlooks the varying information density of speech and, critically, destroys the fine-grained temporal dynamics [stevens2002toward]. This frame-based segmentation arbitrarily cuts through natural speech events, disrupting prosodic rhythms, distorting pause structures, and obscuring speech rate variations. These temporal dynamics are not noise; they are established as primary biomarkers in critical clinical applications, most notably depression detection [huang2019investigation, zhang2025speecht]. Thus, a fundamental mismatch exists: the dominant speech representation methods are optimized for linguistic content, making them inherently unsuitable for downstream tasks that depend on temporal fidelity.
Figure 1: Comparison of segmentation strategies. The Frame-based processing (top) imposes a rigid grid of fixed-length intervals, often disrupting natural speech events. In contrast, Distinctive feature-based processing (bottom) adaptively segments the signal by identifying perceptually significant acoustic transitions as boundaries. This results in variable-length segments that preserve the fine-grained temporal dynamics essential for clinical tasks
To bridge this gap, we must identify meaningful units within continuous signals in a way that preserves temporal structure. The linguistic theory of distinctive features provides an insightful alternative [jakobson1952preliminaries]. This theory posits that speech can be naturally segmented at points where acoustic characteristics are most differentiated [liu1996landmark, zhang2024auto]. Instead of arbitrary fixed-length segments, this approach identifies boundaries where acoustic properties undergo significant changes. Such an adaptive, variable-length segmentation process naturally encodes the temporal dynamics; for example, a pause is no longer a sequence of identical ”silence” frames, but is represented by a single, long segment whose duration is explicitly preserved. However, implementing distinctive feature analysis in neural networks has remained challenging. The irregular, variable-length nature of these features conflicts with the regular, grid-like computations (e.g., convolutions, transformers) that dominate modern deep learning [lecun2015deep, krizhevsky2012imagenet, baevski2020wav2vec, hsu2021hubert]. This technical mismatch has largely confined distinctive feature analysis to traditional signal processing approaches [stevens2002toward, li2008contribution].
In this paper, we address this challenge by introducing the Distinctive Feature Codec (DFC), a framework designed to learn an efficient speech representation that preserves the critical temporal information required for depression detection. We propose the first architecture that successfully integrates the theory of distinctive features into a modern, end-to-end neural codec framework [sutskever2014sequence, hinton2006reducing]. Our approach uses a lightweight, self-supervised boundary detector to identify perceptually significant acoustic transitions, which guide an adaptive encoder-decoder (based on SEANet [tagliasacchi2020seanet]) to process variable-length segments. This marks a departure from conventional fixed-interval processing [defossezhigh, zeghidour2021soundstream, zhang2024speechtokenizer]. Furthermore, our investigation reveals that standard quantization methods like Finite Scalar Quantization (FSQ) [mentzerfinite] become unstable when applied to such variable-length segments at low bitrates. To solve this, we develop a novel Group-wise Scalar Quantization (GSQ) approach, which ensures robust and stable quantization. Our work validates distinctive features as a promising direction for codec design, offering new perspectives on efficient speech representation for depression detection.
II Preliminary
II-A Distinctive Features
@@
Distinctive features, first proposed in linguistic theory [jakobson1952preliminaries, liu1996landmark, lee1988automatic, zhang2024auto], characterize speech by identifying regions with acoustically distinctive properties that help differentiate speech segments from one another. As illustrated in Fig.1, this approach fundamentally differs from conventional frame-based processing: while frame-based methods uniformly segment speech signals into fixed-length overlapping windows, distinctive feature analysis identifies boundaries where acoustic characteristics undergo significant changes. This approach has proven valuable in early automatic speech recognition systems [liu1996landmark, lee1988automatic, he2019ctc] and has been successfully applied to various healthcare applications [ishikawa2017toward, ishikawa2023landmark, zhang2024llms, zhang2025speecht, zhang2025pre]. However, despite its theoretical advantages, the development of distinctive feature-based methods has faced significant limitations in the deep learning era. Traditional implementations of distinctive features heavily rely on linguistic expertise and hand-crafted rules [liu1996landmark, zhang2024auto], leading to limited training data and difficulties in scaling across different acoustic conditions and languages. Additionally, while frame-level processing naturally aligns with convolutional neural networks and enables efficient batch processing, the variable-length nature of distinctive features poses challenges for modern deep learning architectures. The success of frame-level processing in various deep learning systems [baevski2020wav2vec, hsu2021hubert] has led to a rich ecosystem of pre-trained models and established practices, making it the predominant choice for modern speech processing systems despite its inherent inefficiencies. This technical mismatch, combined with the limited availability of labeled data for distinctive feature analysis, has constrained its adoption in contemporary deep learning approaches.
II-B Speech Codecs
Speech codecs compress speech signals into discrete tokens while preserving essential acoustic and linguistic information [vinton2001scalable, epps2000wideband]. These discrete representations serve as inputs for downstream tasks such as large language models [du2024cosyvoice, du2024cosyvoice2, defossez2024moshi]. Current approaches can be broadly categorized into two paradigms: end-to-end trained codecs that directly learn discrete representations through reconstruction objectives [zeghidour2021soundstream, defossezhigh, zhang2024speechtokenizer], and two-stage approaches that first extract semantic features using self-supervised models [hsu2021hubert, baevski2020wav2vec], then apply generative modeling techniques such as flow matching [lipmanflow, mehta2024matcha] or diffusion [ho2020denoising, zhang2023speak].
Two-stage approaches (e.g., Sylber [cho2024sylber]) achieve impressive compression by operating at syllable-level granularity [bpe, baade2024syllablelm], reducing token rates to as low as 5-10 tokens per second. These methods prioritize semantic preservation—retaining linguistic content that overlaps with text representations—while relying on powerful generative models. However, this design philosophy is fundamentally misaligned with depression detection, where diagnostic information resides not in semantic content but in acoustic characteristics: voice quality, spectral energy distribution, and prosodic micro-variations [zhang2025speecht].
Table I, based on our experimental evaluation of the Sylber model, empirically demonstrates this limitation. As Sylber’s model capacity increases (K-means clusters from 10K to 20K), perceptual quality (PESQ) paradoxically degrades from 1.25 to 0.81, while mel-spectral error increases from 0.616 to 0.673. This pattern indicates that the generative model produces perceptually plausible speech by hallucinating acoustic details rather than preserving them from the original signal. While the semantic content (”what was said”) remains intact, the acoustic substrate (”how it was said”) diverges from the input—erasing the subtle spectral and prosodic deviations that constitute depression biomarkers. Multiple studies confirm that such semantically-focused representations underperform on tasks requiring acoustic fidelity [cuervo2024scaling, wang2024speech].
TABLE I: Analysis of Sylber Codec performance. The table shows mel_error, stft, pseq, and stoi metrics for different KM (K-Means) cluster sizes.
Sylber Codec mel_error $\downarrow$ stft $\downarrow$ PESQ $\uparrow$ stoi $\uparrow$
KM=5K 0.6204 2.0246 1.0025 0.7066
KM=10K 0.6155 2.2049 1.2501 0.7136
KM=20K 0.6727 2.1525 0.8135 0.7003
Figure 2: Overview of the Distinctive Feature Codec framework. The Distinctive Feature Detector (top) identifies acoustic boundaries through contrastive learning to guide variable-length segmentation. The codec pipeline segments encoded speech based on these boundaries, compresses each segment via the DF Encoder, applies quantization, and reconstructs through the DF Decoder and main decoder.
Our work therefore focuses on end-to-end trained codecs, which optimize for faithful acoustic reconstruction rather than semantic compression, explicitly preserving the spectral and temporal dynamics essential for clinical applications.
II-C Speech Representation for Depression Detection
Clinical research has long established that depression profoundly affects speech production mechanisms, manifesting in distinctive acoustic biomarkers known as “psychomotor retardation” [sobin1997psychomotor]. These manifestations primarily include reduced speech rate, monotonic prosody, prolonged pauses, and blunted articulation [mundt2012vocal, quatieri2012vocal]. Consequently, capturing these temporal dynamics and prosodic variations has been a central focus in feature extraction for depression detection.
Early approaches predominantly relied on hand-crafted Low-Level Descriptors (LLDs) to explicitly model these characteristics. Standard feature sets, such as eGeMAPS and INTERSPEECH ComParE [eyben2015geneva], aggregate frame-level acoustic properties (e.g., pitch, energy, formants) using statistical functionals to capture global prosodic trends. More distinctively, some studies explored landmark-based or distinctive feature analysis [huang2019investigation, zhang2025pre], which focuses on specific acoustic events—such as the onset of bursts or glottal transitions—to detect subtle articulatory coordination deficits associated with depression [huang2019investigation, ishikawa2017toward, zhang2025speecht]. These methods offered high interpretability and preserved the temporal integrity of speech events but were limited by their inability to model complex, high-level abstractions.
In the deep learning era, the paradigm shifted towards learning data-driven representations directly from raw audio or spectrograms using Convolutional Neural Networks (CNNs) and Transformers [lecun2015deep, krizhevsky2012imagenet]. Self-supervised models (SSL) like Wav2Vec 2.0 and HuBERT have achieved state-of-the-art performance by encoding speech into continuous or discrete representations [baevski2020wav2vec, hsu2021hubert, zeghidour2021soundstream, zhang2024speechtokenizer]. However, these modern approaches predominantly employ fixed-rate frame-based processing, where speech is segmented into uniform time intervals regardless of the underlying acoustic content. While effective for linguistic tasks like ASR, recent studies suggest that this rigid segmentation may disrupt the fine-grained temporal structures—such as variable-length pauses and rhythmic patterns—that serve as critical diagnostic cues [stevens2002toward, zhang2025pre, zhang2025speecht]. This creates a fundamental mismatch: the dominant feature extraction methods are optimized for semantic continuity, potentially at the cost of the temporal fidelity required for reliable depression assessment.
III Distinctive Codec Framework
@@
As illustrated in Fig. 2, our approach consists of two key stages: First, a lightweight boundary detector is trained to identify perceptually significant transitions in speech signals. Second, these detected boundaries guide the codec model to adaptively merge or separate speech segments, leading to more efficient tokenization that aligns with the natural structure of speech.
III-A Distinctive Features Detector
@@
The core idea of distinctive features lies in identifying regions where speech segments exhibit maximal acoustic contrast with their neighbors. This naturally aligns with the objective of contrastive learning, which aims to learn representations by maximizing the differences between distinct samples while minimizing differences between similar ones [khosla2020supervised, gao2021simcse, reimers2019sentence, kreuk2020self]. We leverage this connection to design a self-supervised boundary detector that learns to identify distinctive features without requiring phoneme-level annotations.
@@
III-B Distinctive Codec
To evaluate our distinctive feature-based approach, we build our codec model on top of the SpeechTokenizer framework [zhang2024speechtokenizer], which leverages the SEANet architecture [tagliasacchi2020seanet] for encoder-decoder operations. The key innovation of our approach lies in how we process the encoded representations based on the distinctive features detected by our boundary detector. Given input speech $\mathbf{x}\in\mathbb{R}^{1\times L}$, the encoder $E_{\theta}$ first maps it to a latent representation:
@@ III-B Distinctive Codec
where $D$ is the feature dimension and $T=L/r$ represents the temporal dimension after downsampling with ratio $r$. Traditional frame-based codecs would typically process this representation uniformly across time. In contrast, our distinctive codec first identifies segment boundaries $\mathcal{B}={b_{1},b_{2},\ldots,b_{M}}$ using the boundary detector described in Section III-A. With these boundaries, we partition the feature sequence into variable-length segments:
@@
where $\mathbf{s}_{i}=\mathbf{e}[:,b_{i-1}:b_{i}]$ represents the $i$-th segment ($b_{0}=0$ and $b_{M+1}=T$ for notational convenience). For each segment, we apply a Distinctive Feature encoder (DFE) to compress the variable-length representation into a fixed-length embedding:
@@
[…]
III-C Group-wise Scalar Quantization
Finite Scalar Quantization (FSQ) [mentzerfinite] has emerged as an effective approach for discrete representation learning due to its computational efficiency and strong performance across various tasks [du2024cosyvoice2, yu2024language, parker2024scaling]. Unlike vector quantizers that require nearest neighbor search in high-dimensional spaces, FSQ directly quantizes each dimension of the latent representation independently, significantly reducing computational complexity.
@@ III-C Group-wise Scalar Quantization
However, during our experiments with distinctive feature-based tokenization, we discovered that standard FSQ becomes unstable when operating at high downsampling rates. This instability manifests as training divergence and poor reconstruction quality, particularly when compressing longer, variable-length segments into single tokens. We hypothesize that this issue stems from the increased difficulty of directly quantizing high-dimensional features with varying temporal characteristics.
To address this issue, we propose Group-wise Scalar Quantization (GSQ) as shown in Figure 3, which decomposes the high-dimensional quantization problem into multiple lower-dimensional sub-problems. Given a compressed segment representation $\mathbf{z}_{i}\in\mathbb{R}^{H\times 1}$, we divide it into $G$ groups, each with dimension $H_{g}=H/G$:
$\mathbf{z}_{i}=[\mathbf{z}_{i}^{1},\mathbf{z}_{i}^{2},\ldots,\mathbf{z}_{i}^{G}]$ (9)
Each group is then processed independently through quantization. The key design choice lies in how we parameterize the quantization function for each group $Q_{g}(\cdot)$. We explore two variants:
Direct Quantization (M2M): Each group is quantized directly without dimensional change:
$\hat{\mathbf{z}}_{i}^{g}=\text{FSQ}_{g}(\mathbf{z}_{i}^{g}),\quad\mathbf{z}_{i}^{g},\hat{\mathbf{z}}_{i}^{g}\in\mathbb{R}^{H_{g}}$ (10)
This preserves the distributed representation across all $H_{g}$ dimensions but requires a codebook of size $L^{H_{g}}$ per group (where $L$ is the number of quantization levels per dimension).
Projection-based Quantization (M2O): Each group is first projected to a lower-dimensional space (in our case, a scalar), quantized, then projected back:
$p_{i}^{g}=\mathbf{W}_{g}^{(c)}\mathbf{z}_{i}^{g},\quad\hat{p}_{i}^{g}=\text{FSQ}_{g}(p_{i}^{g}),\quad\hat{\mathbf{z}}_{i}^{g}=\mathbf{V}_{g}^{(e)}\hat{p}_{i}^{g}$ (11)
where $\mathbf{W}_{g}^{(c)}\in\mathbb{R}^{k\times H_{g}}$ is the compression matrix, $\mathbf{V}_{g}^{(e)}\in\mathbb{R}^{H_{g}\times k}$ is the expansion matrix, and $k$ is the projection dimension (we use $k=1$, i.e., scalar projection). This reduces the codebook size to $L^{k}$ per group while maintaining expressiveness through learned projections.
@@
$\hat{\mathbf{z}}_{i}=[\hat{\mathbf{z}}_{i}^{1},\hat{\mathbf{z}}_{i}^{2},\ldots,\hat{\mathbf{z}}_{i}^{G}]$ (12)
This group-wise decomposition is particularly beneficial for quantizing variable-length segments with diverse acoustic characteristics. By processing features in smaller groups with dedicated quantizers, we reduce quantization complexity through lower-dimensional operations. Our primary experiments employ the M2O variant with scalar projection ($k=1$) due to its superior stability at low token rates and dramatically reduced codebook size. Algorithm 1 presents the unified framework, with the projection step (lines 3-4, 6) being optional for the M2M variant.
Algorithm 1 Group-wise Scalar Quantization (GSQ)
0: Input tensor $x\in\mathbb{R}^{B\times T\times D}$
0: Number of groups $G$, group size $d=D/G$
0: Use projection: $\texttt{use\_proj}\in\{\text{true},\text{false}\}$
0: Projection dimension: $k$ (typically $k=1$ for M2O)
0:If use_proj: compression matrices $\{W_{i}^{(c)}\in\mathbb{R}^{k\times d}\}_{i=1}^{G}$, expansion matrices $\{W_{i}^{(e)}\in\mathbb{R}^{d\times k}\}_{i=1}^{G}$
0: FSQ quantizers $\{\text{FSQ}_{i}\}_{i=1}^{G}$
0: Quantized output $\hat{x}\in\mathbb{R}^{B\times T\times D}$
1:
2:for $i=1$ to $G$ do
3:  $x_{i}\leftarrow x[:,:,i\cdot d:(i+1)\cdot d]$ {Extract group ii}
4:  if use_proj then {M2O: project to lower dim}
5:   $x_{i}\leftarrow W_{i}^{(c)}x_{i}$
6:  end if
7:  $\hat{x}_{i}\leftarrow\text{FSQ}_{i}(x_{i})$ {Quantize (M2M or M2O)}
8:  if use_proj then {M2O: project back}
9:   $\hat{x}_{i}\leftarrow W_{i}^{(e)}\hat{x}_{i}$
10:  end if
11:end for
12: